OpenStudy (gabbyalicorn):

The area of the base of a prism is 21 cm2. The perimeter of the base is 20 cm. The height of the prism is 8 cm. What is the surface area of the prism?

OpenStudy (anonymous):

what is the shape of prism?

OpenStudy (gabbyalicorn):

there is no shape that they gave me

OpenStudy (anonymous):

let it be rectangular and x and y be its sides. 2(x+y)=20 x+y=20/2=10 y=10-x xy=21 x(10-x)=21 10x-x^2=21 x^2-10x+21=0 x^2-7x-3x+21=0 x(x-7)-3(x-7)=0 (x-7)(x-3)=0 either x-7=0,x=7,y=10-7=3 or x-3=0,x=3,y=10-3=7 so length =7 cm breadth=3 cm surface area=2*21+20*8=? cm

OpenStudy (gabbyalicorn):

496

OpenStudy (anonymous):

42+160=?